# Solutions of the singular Yamabe problem near singular boundaries

### v1

@@ 2. Preliminaries
for any $x\in\Omega\cap B_{r_{1}}(0)$, by taking $r_{1}$ smaller.
@@ 5. Examples with no tangential decay
In this section, we will introduce some examples where the tangential decay is absent. Specifically, let $\delta>0$ be a constant, we will demonstrate some examples, where the estimate in Theorem 1.2 can not be improved as the following:
(5.1) $\left|\frac{u(x)}{u_{V}(Tx)}-1\right|\leq Cd_{g}(x,0)\rho^{\delta}(\theta)$
for $x\in\Omega\cap B_{r_{0}}$ and $\theta=Tx/|Tx|$. This constitutes the key difference from the results in [HJS]. In [HJS], where the domains are cones, the aforementioned estimate holds well, and it is instrumental in deriving higher-order expansions.
We begin by reviewing solutions of the Loewner-Nirenberg problem (1.3)-(1.4) in cones that possess translational invariance. As a preparatory step for our first example, we focus on the specific scenario with $n=3$, and the invariance is along the $z$-axis.
@@
Let $V_{\alpha}$ be a wedge in $\operatorname{\mathbb{R}}^{2}$ with the origin as its vertex, and with the opening angle $\alpha\in(0,\pi)$. Without loss of generality, we assume that $\partial V_{\alpha}=\{\theta=0\}\cup\{\theta=\alpha\}$, where $\theta\in(0,\pi)$ is defined by $(\cos\theta,\sin\theta)=\frac{1}{\sqrt{x^{2}+y^{2}}}(x,y)$. Suppose $V=V_{\alpha}\times\operatorname{\mathbb{R}}:=\{(x,y,z):(x,y)\in V_{\alpha}\}\subset\operatorname{\mathbb{R}}^{3}$, and $u_{V}(x,y,z)$ is the positive solution to the Loewner-Nirenberg problem (1.3)-(1.4) in $V$. Then,
$u_{V}(x,y,z)=|\sqrt{x^{2}+y^{2}}|^{-\frac{1}{2}}\eta(\theta),$
where $\theta\in(0,\alpha)$, and $\eta\in C^{\infty}(0,\alpha)$ satisfies
(5.2) $\left\{\begin{aligned} &\eta^{\prime\prime}(\theta)+\frac{1}{4}\eta(\theta)=\frac{3}{4}\eta^{5}(\theta)\quad\mbox{ in }(0,\alpha),\\ &\eta(0)=+\infty,\quad\eta(\alpha)=+\infty.\end{aligned}\right$.
The proof of this proposition can be found in [HS]. Recall that in Section 1 we also present another form of the solution of the Loewner–Nirenberg problem (1.3)–(1.4) in cones. By the uniqueness, this represents the same solution expressed in a different form.
@@
$\Omega=\left\{(x,y,z)\in\operatorname{\mathbb{R}}^{3}:0\leq y\leq\frac{1+z}{100}x\right\}.$
Suppose $V_{z}$ is the tangent cone of $\Omega$ at $(0,0,z)$, then $V_{z}$ is translationally invariant along the $z$-axis. For $i=1,2$, let $P_{i}^{z}$ be the tangent plane of $S_{i}$ at $(0,0,z).$ As in [HS], for $q\in V_{z}$, we define $f_{V_{z}}(d_{1},d_{2})$ by letting $f_{V_{z}}(d_{P_{1}^{z}}(q),d_{P_{2}^{z}}(q))=u_{V_{z}}(q)$ be the solution of the Loewner-Nirenberg problem (1.3)-(1.4) in $V_{z}$, where $d_{P_{i}^{z}}(q)$ is the distance from $q$ to $P^{z}_{i}$.
Suppose $u>0$ solves the Loewner-Nirenberg problem (1.3)-(1.4) in $\Omega$. Let $z>0$ and $p=(x,y,z)\in\Omega$ be a point such that the distances from $p$ to $S_{1}$ ans $S_{2}$ are $ks$ and $s$ respectively, where $s>0$ and $k\in(0,1)$ are positive constants. When $p$ is close to the origin, by [HS] we know
@@
By Proposition 5.1 we know $\frac{f_{V_{z}}(ks,s)}{f_{V_{0}}(ks,s)}$ is independent of $s$. We claim there are some $z>0$ and $k\in(0,1)$ such that
@@
Next, we present a more general method for constructing examples that exhibit no tangential decay. Surprisingly, even when the boundary has higher regularity and its tangent cone is a half-space, tangential decay is not guaranteed.
@@
where $u_{H}=r^{-\frac{n-2}{2}}\xi$ is the solution of the Loewner-Nirenberg problem in $\operatorname{\mathbb{R}}^{n}_{+}$, and $c_{1}(\theta)=\frac{\xi_{1}(\theta)}{\xi(\theta)}$. Here, $\xi$ is the solution of (1.5)-(1.6) in $\Sigma:=\partial B_{1}\cap\operatorname{\mathbb{R}}_{+}^{n}$, and $\xi_{1}$ is a solution of (4.1) in $\Sigma$ with
@@
$a_{ij}=(g^{kl}T^{i}_{k}T^{j}_{l})(Sy)=\delta_{ij}+a_{ij,k}y_{k}+O(r^{2}),$
$b_{i}=\left(g^{kl}T_{kl}^{i}+\frac{1}{\sqrt{\mathrm{det}g}}\partial_{x_{k}}(\sqrt{\mathrm{det}g}g^{kl})T^{i}_{l}\right)(Sy)=b_{i,0}+O(r),$
$g^{ij}=\delta_{ij}+O(r^{2}).$
@@
Note $r^{\frac{n}{2}}b_{i,0}(u_{H})_{i}=O(\xi^{\frac{n}{n-2}})$, and $(y_{1}+\cdots+y_{n})r^{-1}\xi^{\frac{n+2}{n-2}}\sim\xi^{\frac{n+2}{n-2}},$ we know $F\sim\xi^{\frac{n+2}{n-2}}$, and hence $\rho^{2}F\sim\xi$. Note $\mu_{1}>1$ (by Remark 1.4), if (5.1) holds, then $\xi_{1}=O(\xi^{1-\frac{2\delta}{n-2}})$, which is impossible when we compare the order of two sides of the equation (4.1). Therefore, (5.1) dose not hold in this case.
@@ Appendix A Existence of Solutions to Degenerate/Singular Elliptic Equations
Let $\Sigma\subset\mathbb{S}^{n-1}$ be a Lipschitz domain, $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6),
@@
and $d_{\Sigma}=d_{g_{\mathbb{S}^{n-1}}}(\cdot,\partial\Sigma).$ By Lemma 2.4 in [HJS], we have $\rho\in C^{\infty}(\Sigma)\cap\mathrm{Lip}(\bar{\Sigma})$, $\rho>0$ in $\Sigma$, $\rho=0$ on $\partial\Sigma$, and $\rho\sim d_{\Sigma}$. Denote by $\{\theta_{i}\}^{n-1}_{i=1}$ normal local coordinates on the sphere $\mathbb{S}^{n-1}$. Set $\partial_{i}=\partial_{\theta_{i}}$ for $1\leq i\leq n-1$, and $\partial_{ij}=\partial_{i}\partial_{j}=\partial_{j}\partial_{i}$. We define the uniformly degenerate elliptic operator $L_{0}$ on $\Sigma$ by
@@
where $a_{ij},b_{i},c\in L^{\infty}(\Sigma)$ with
(A.3) $\lambda|\eta|^{2}\leq a_{ij}\eta_{i}\eta_{j}\leq\Lambda|\eta|^{2},$
@@
We now present a proof of the existence result for uniformly degenerate elliptic equations in Lipschitz domains, due to Graham and Lee [GL]. A proof is provided here for the sake of completeness.
@@
Let $\Sigma\subset\mathbb{S}^{n-1}$ be a Lipschitz domain, $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6), and $a_{ij},b_{i},c\in L^{\infty}(\Sigma)\cap C^{\alpha}(\Sigma)$ with (A.3), for some $\alpha\in(0,1)$. Assume there exists a function $\psi\in C^{2}(\Sigma)$ such that $\psi>0$ in $\Sigma$, $\rho\nabla_{\theta}\log\psi\in L^{\infty}(\Sigma)$, and, for some constant $\delta>0$,
@@
Consider a sequence of increasing domains $\{\Omega_{k}\}^{\infty}_{k=1}$ with $C^{2,\alpha}$-boundaries such that $\Omega_{k}\subset\subset\Omega_{k+1}\subset\subset\Sigma$ and $\Sigma=\cup^{\infty}_{k=1}\Omega_{k}$. Since $\rho>0$ in $\Sigma$, the operator $L_{0}$ is uniformly elliptic in $\Omega_{k}$ with $C^{\alpha}(\bar{\Omega}_{k})$-coefficients. By the Schauder theory, there exists a unique solution $u_{k}\in C^{2,\alpha}(\bar{\Omega}_{k})$ of
@@
Set $v_{k}=u_{k}/\psi$. A direct computation yields $v_{k}$ satisfies
@@
By applying the interior Schauder estimates to $L_{0}u_{k}=f$ in $\Omega_{i}\subset\subset\Omega_{i+1}$, we obtain, for any $k\geq i+1$,
@@
$\leq C\Big\{\frac{1}{\delta}\|\psi\|_{L^{\infty}(\Omega_{i+1})}\Big\|\frac{f}{\psi}\Big\|_{L^{\infty}(\Sigma)}+\|f\|_{C^{\alpha}(\bar{\Omega}_{i+1})}\Big\}$
where $C$ is a positive constant depending on $i$, but independent of $k$. By a standard diagonalization, there exist a subsequence $\{u_{k^{\prime}}\}$ of $\{u_{k}\}$ and a function $u\in C^{2,\alpha}(\Sigma)$ such that $u_{k^{\prime}}\rightarrow u$ in $C^{2}(\bar{\Omega}_{i})$ for any $i\geq 1$. Thus, $L_{0}u=f$ in $\Sigma$. By taking $k\rightarrow\infty$ in (A.7), we obtain (A.6). ∎
@@
(A.8) $L_{1}=\Delta_{\theta}-\frac{\kappa}{\rho^{2}}\quad\text{on }\Sigma,$
@@
Here we collect some useful results from [HJS, SW].
@@
Let $\Sigma\subset\mathbb{S}^{n-1}$ be a Lipschitz domain and $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6).
$\mathrm{(1)}$(([HJS], Theorem 4.2)) There exists an increasing sequence of positive constants $\{\lambda_{i}\}_{i\geq 1}$, divergent to $\infty$, and an $L^{2}(\Sigma)$-orthonormal basis $\{\phi_{i}\}_{i\geq 1}$ such that, for $i\geq 1$, $\phi_{i}\in C^{\infty}(\Sigma)\cap H^{1}_{0}(\Sigma)$ and $L_{1}\phi_{i}=-\lambda_{i}\phi_{i}$ weakly.
$\mathrm{(2)}$(([HJS], Theorem 4.3)) For any $\lambda\notin\{\lambda_{i}\}$ and any $f\in L^{2}(\Sigma)$, there exists a unique weak solution $u\in H^{1}_{0}(\Sigma)$ of the equation
(A.9) $L_{1}u+\lambda u=f\quad\text{in }\Sigma.$
Moreover,
@@
where $C$ is a positive constant depending only on $n$, $\lambda$, and $\Sigma$.
$\mathrm{(3)}$(([SW], Proposition 4.5)) If $n=3$, then $\lambda_{1}>\frac{3}{4}$.
$\mathrm{(4)}$(([HJS], Theorem 4.4; [SW], Lemma 4.1)) Assume that for some $\lambda\in\mathbb{R}$ and $f\in C^{\infty}(\Sigma)$, $u\in H^{1}_{0}(\Sigma)$ is a weak solution of (A.9). Then, there exists a constant $\nu>0$ depending only on $n$ and $\Sigma$ such that if, for some constant $A>0$,
@@
where $C$ is a positive constant depending only on $n$, $\lambda$, and $\Sigma$. In particular, let $\phi_{i}$ be an eigenfunction as in $\mathrm{(1)}$. Then,
(A.12) $|\phi_{i}|+\rho|\nabla_{\theta}\phi_{i}|+\rho^{2}|\nabla^{2}_{\theta}\phi_{i}|\leq C\rho^{\nu}\quad\text{in }\Sigma.$
where C is a positive constant depending only on $n$, $\lambda_{i}$, and $\Sigma$.

### v2

@@ 2. Preliminaries
for any $x\in\Omega\cap B_{r_{1}}(0)$, by taking $r_{1}$ smaller. ∎
Remark 4.3.
When $T$ is only $C^{2,\alpha}$ with some $\alpha\in(0,1]$, we can still obtain the first order expansion of $u/(u_{V}\circ T)$. The proof strategies are similar. In fact, we can construct the related function $w$ as follows.
  1. (1)
    $w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{1+\alpha}+A_{1}\phi_{1}r^{-\frac{n}{2}+2+\alpha}$, if $\mu_{1}>1+\alpha$;
  2. (2)
    $w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{1+\alpha}-A_{1}\phi_{1}r^{-\frac{n}{2}+2+\alpha}\log r$, if $\mu_{1}=1+\alpha$;
  3. (3)
    $w=u_{V}+\xi_{1}r^{-\frac{n}{2}+2}+A_{0}u_{V}r^{1+\alpha}+A_{1}\phi_{1}(r^{\mu_{1}-\frac{n-2}{2}}-r^{-\frac{n}{2}+2+\alpha})$, if $\mu_{1}<1+\alpha$.
A similar analysis, combined with the inductive argument, leads to the result in Remark 1.5.
@@ 5. Examples with no tangential decay
In this section, we will introduce some examples in which the tangential decay is absent. Specifically, we will demonstrate some examples, where the estimate in Theorem 1.2 can not be improved as the following:
(5.1) $\left|\frac{u(x)}{u_{V}(Tx)}-1\right|\leq Cd_{g}(x,0)o(1),$
for $x\in\Omega\cap B_{r_{0}}(0)$ and $\theta=Tx/{|Tx|}$, where $o(1)\to 0$ as $\rho(\theta)\to 0$ and $\rho\sim d_{\Sigma}$ in $\Sigma$. This constitutes the key difference from the results in [10]. In [10], where the domains are cones, the aforementioned estimate holds well and is instrumental in deriving higher-order expansions.
We begin by reviewing solutions of the Loewner-Nirenberg problem (1.3)-(1.4) in cones with translational invariance. As a preparatory step for our first example, we focus on the specific scenario with $n=3$, and the invariance is along the $z$-axis.
@@
Let $V_{\alpha}$ be a wedge in $\operatorname{\mathbb{R}}^{2}$ with the origin as its vertex and with the opening angle $\alpha\in(0,\pi)$. Assume that $\partial V_{\alpha}=\{\tau=0\}\cup\{\tau=\alpha\}$, where $\tau\in(0,\pi)$ is defined by $(\cos\tau,\sin\tau)=\frac{1}{\sqrt{x^{2}+y^{2}}}(x,y)$. Suppose $V=V_{\alpha}\times\operatorname{\mathbb{R}}:=\{(x,y,z):(x,y)\in V_{\alpha},z\in\mathbb{R}\}\subset\operatorname{\mathbb{R}}^{3}$, and $u_{V}(x,y,z)$ is the positive solution to the Loewner-Nirenberg problem (1.3)-(1.4) in $V$. Then
$u_{V}(x,y,z)=|\sqrt{x^{2}+y^{2}}|^{-\frac{1}{2}}\eta(\tau),$
where $\tau\in(0,\alpha)$, and $\eta\in C^{\infty}(0,\alpha)$ satisfies
(5.2) $\left\{\begin{aligned} &\eta^{\prime\prime}(\tau)+\frac{1}{4}\eta(\tau)=\frac{3}{4}\eta^{5}(\tau)\quad\mbox{ in }(0,\alpha),\\ &\eta(0)=+\infty,\quad\eta(\alpha)=+\infty.\end{aligned}\right$.
The proof of this proposition can be found in [12]. Recall that in Section 1 we also present another form of the solution of the Loewner–Nirenberg problem (1.3)–(1.4) in cones. By the uniqueness, this represents the same solution expressed in a different form.
@@
$\Omega=\left\{(x,y,z)\in B_{1}(0):0\leq y\leq\frac{1+z}{100}x\right\}.$
Suppose $V_{z}$ is the tangent cone of $\Omega$ at $(0,0,z)$, then $V_{z}$ is translationally invariant along the $z$-axis. For $i=1,2$, let $P_{i}^{z}$ be the tangent plane of $S_{i}$ at $(0,0,z).$ As in [12], for $q\in V_{z}$, we define $f_{V_{z}}(d_{1},d_{2})$ by letting $f_{V_{z}}(d_{P_{1}^{z}}(q),d_{P_{2}^{z}}(q))=u_{V_{z}}(q)$ be the solution of the Loewner-Nirenberg problem (1.3)-(1.4) in $V_{z}$, where $d_{P_{i}^{z}}(q)$ is the distance from $q$ to $P^{z}_{i}$.
Suppose $u>0$ solves the Loewner-Nirenberg problem (1.3)-(1.4) in $\Omega$. Let $z>0$ and $p=(x,y,z)\in\Omega$ be a point such that the distances from $p$ to $S_{1}$ ans $S_{2}$ are $ks$ and $s$ respectively, where $s>0$ and $k\in(0,1)$ are positive constants. When $p$ is close to the origin, by [12] we know
@@
By Proposition 5.1, we know $\frac{f_{V_{z}}(ks,s)}{f_{V_{0}}(ks,s)}$ is independent of $s$. We claim there are some $z>0$ and $k\in(0,1)$ such that
@@
Next, we present a more general method for constructing examples that exhibit no tangential decay. Surprisingly, even when the boundary is smooth and its tangent cone is a half-space, tangential decay is not guaranteed.
@@
where $u_{H}=r^{-\frac{n-2}{2}}\xi$ is the solution of the Loewner-Nirenberg problem in $\operatorname{\mathbb{R}}^{n}_{+}$, and $c_{1}(\theta)=\frac{\xi_{1}(\theta)}{\xi(\theta)}$. Here, $\xi$ is the solution of (1.5)-(1.6) in $\Sigma:=\partial B_{1}\cap\operatorname{\mathbb{R}}_{+}^{n}$, and $\xi_{1}$ is a solution of (4.1) in $\Sigma$ with $k=1$, $\sigma=0$,
@@
$a_{ij}=(g^{kl}T^{i}_{k}T^{j}_{l})(Sy)=\delta_{ij}+a_{ij,k}y_{k}+O(r^{2}),\mbox{ as }r\to 0,$
$b_{i}=\left(g^{kl}T_{kl}^{i}+\frac{1}{\sqrt{\mathrm{det}g}}\partial_{x_{k}}(\sqrt{\mathrm{det}g}g^{kl})T^{i}_{l}\right)(Sy)=b_{i,0}+O(r),\mbox{ as }r\to 0,$
$g^{ij}=\delta_{ij}+O(r^{2}),\mbox{ as }r\to 0.$
@@
Let $Q:=\partial B_{1}\cap\{y\in\operatorname{\mathbb{R}}^{n}_{+}:y_{1}>0,y_{2}>0,\cdots,y_{n-1}>0\}\subset\Sigma$, then, $r^{\frac{n}{2}}b_{i,0}(u_{H})_{i}\lesssim\xi^{\frac{n}{n-2}}$, and $(y_{1}+\cdots+y_{n})r^{-1}\xi^{\frac{n+2}{n-2}}\sim\xi^{\frac{n+2}{n-2}}$ in $Q$. Hence, we know $F\sim\xi^{\frac{n+2}{n-2}}$, and further $\rho^{2}F\sim\xi$ in $Q$. Note $\mu_{1}>1$ (by Remark 1.4), if (5.1) holds, then $\xi_{1}=o(1)\xi$ in $Q$, where $o(1)\to 0$ as $\rho(\theta)\to 0$. It is impossible when we compare the order of two sides of the equation (4.1) in $Q$. Therefore, (5.1) dose not hold in this case.
@@ Appendix A Existence of Solutions to Degenerate/Singular Elliptic Equations
Let $\Sigma\subsetneq\mathbb{S}^{n-1}$ be a Lipschitz domain, $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6),
@@
and $d_{\Sigma}=d_{g_{\mathbb{S}^{n-1}}}(\cdot,\partial\Sigma).$ By Lemma 2.4 in [10], we have $\rho\in C^{\infty}(\Sigma)\cap\mathrm{Lip}(\bar{\Sigma})$, $\rho>0$ in $\Sigma$, $\rho=0$ on $\partial\Sigma$, and $\rho\sim d_{\Sigma}$. Denote by $\{\theta_{i}\}^{n-1}_{i=1}$ normal local coordinates on the sphere $\mathbb{S}^{n-1}$. Set $\partial_{i}=\partial_{\theta_{i}}$ and $\partial_{ij}=\partial_{i}\partial_{j}=\partial_{j}\partial_{i}$ for $i,j=1,\cdots,n-1$. We define the uniformly degenerate elliptic operator $L_{0}$ in $\Sigma$ by
@@
where $a_{ij},b_{i},c\in L^{\infty}(\Sigma)$ satisfying $a_{ij}=a_{ji}$, and for any $\theta\in\Sigma$ and $\eta\in\mathbb{R}^{n-1}$,
(A.3) $\lambda|\eta|^{2}\leq a_{ij}(\theta)\eta_{i}\eta_{j}\leq\Lambda|\eta|^{2},$
@@
We now present an existence result for uniformly degenerate elliptic equations in Lipschitz domains, due to Graham and Lee [7]. A proof is provided here for the sake of completeness.
@@
Let $\Sigma\subsetneq\mathbb{S}^{n-1}$ be a Lipschitz domain, $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6), and $a_{ij},b_{i},c\in L^{\infty}(\Sigma)\cap C^{\alpha}(\Sigma)$ with (A.3), for some $\alpha\in(0,1)$. Assume there exists a function $\psi\in C^{2}(\Sigma)$ such that $\psi>0$ in $\Sigma$, $\rho\nabla_{\theta}\log\psi\in L^{\infty}(\Sigma)$, and, for some constant $\delta>0$,
@@
Consider an increasing sequence of domains $\{\Omega_{k}\}^{\infty}_{k=1}$ with boundaries of class $C^{2,\alpha}$ such that $\Omega_{k}\subset\subset\Omega_{k+1}\subset\subset\Sigma$ and $\Sigma=\cup^{\infty}_{k=1}\Omega_{k}$. Since $\rho>0$ in $\Sigma$, the operator $L_{0}$ is uniformly elliptic in $\Omega_{k}$ with coefficients in $C^{\alpha}(\bar{\Omega}_{k})$. By the Schauder theory, there exists a unique solution $u_{k}\in C^{2,\alpha}(\bar{\Omega}_{k})$ of
@@
Define $v_{k}=u_{k}/\psi$. A direct calculation shows that $v_{k}$ satisfies
@@
By applying the interior Schauder estimates to $L_{0}u_{k}=f$ in $\Omega_{i}\subset\subset\Omega_{i+1}$, we get, for any $k\geq i+1$,
@@
$\leq C\Big\{\frac{1}{\delta}\|\psi\|_{L^{\infty}(\Omega_{i+1})}\Big\|\frac{f}{\psi}\Big\|_{L^{\infty}(\Sigma)}+\|f\|_{C^{\alpha}(\bar{\Omega}_{i+1})}\Big\},$
where $C$ is a positive constant depending on $i$, but independent of $k$. Using a standard diagonal argument, we can extract a subsequence $\{u_{k^{\prime}}\}$ of $\{u_{k}\}$ and find a function $u\in C^{2,\alpha}(\Sigma)$ such that $u_{k^{\prime}}\rightarrow u$ in $C^{2}(\bar{\Omega}_{i})$ for any $i\geq 1$. Thus, $L_{0}u=f$ in $\Sigma$. Passing to the limit $k\rightarrow\infty$ in (A.7) yields (A.6). ∎
@@
(A.8) $L_{1}=\Delta_{\theta}-\frac{\kappa}{\rho^{2}}\quad\text{in }\Sigma,$
@@
Here we collect some useful results from [10, 20].
@@
Let $\Sigma\subsetneq\mathbb{S}^{n-1}$ be a Lipschitz domain and let $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6).
$\mathrm{(1)}$(([10, Theorem 4.2])) There exists an increasing sequence of positive constants $\{\lambda_{i}\}_{i\geq 1}$, diverging to $\infty$, and an $L^{2}(\Sigma)$-orthonormal basis $\{\phi_{i}\}_{i\geq 1}$ such that, for $i\geq 1$, $\phi_{i}\in C^{\infty}(\Sigma)\cap H^{1}_{0}(\Sigma)$ and $L_{1}\phi_{i}=-\lambda_{i}\phi_{i}$ holds in the weak sense.
$\mathrm{(2)}$((Fredholm alternative, [10, Theorem 4.3])) For any $\lambda\notin\{\lambda_{i}\}$ and any $f\in L^{2}(\Sigma)$, there exists a unique weak solution $u\in H^{1}_{0}(\Sigma)$ of
(A.9) $L_{1}u+\lambda u=f\quad\text{in }\Sigma,$
satisfying the estimate
@@
where $C>0$ depends only on $n$, $\lambda$, and $\Sigma$.
For any $\lambda=\lambda_{i}$ for some $i$ and any $f\in L^{2}(\Sigma)$ with $(f,\phi_{k})_{L^{2}(\Sigma)}=0$ for any eigenfunction $\phi_{k}$ corresponding to the eigenvalue $\lambda_{i}$, there exists a unique weak solution $u\in H^{1}_{0}(\Sigma)$ of (A.9) such that $(u,\phi_{k})_{L^{2}(\Sigma)}=0$ for any eigenfunction $\phi_{k}$ corresponding to the eigenvalue $\lambda_{i}$, and (A.10) holds.
$\mathrm{(3)}$(([20, Proposition 4.5])) If $n=3$, then $\lambda_{1}>\frac{3}{4}$.
$\mathrm{(4)}$(([10, Theorem 4.4]; [20, Lemma 4.1])) Assume that for some $\lambda\in\mathbb{R}$ and $f\in C^{\infty}(\Sigma)$, $u\in H^{1}_{0}(\Sigma)$ is a weak solution of (A.9). Then, there exists a constant $\nu>0$ depending only on $n$ and $\Sigma$ such that if, for some constant $A>0$,
@@
where $C>0$ depends only on $n$, $\lambda$, and $\Sigma$. In particular, let $\phi_{i}$ be an eigenfunction as in $\mathrm{(1)}$. Then, for any $\beta\in(0,1)$,
(A.12) $|\phi_{i}|+\rho|\nabla_{\theta}\phi_{i}|+\rho^{2}|\nabla^{2}_{\theta}\phi_{i}|+\rho^{2+\beta}[\nabla^{2}_{\theta}\phi_{i}]_{C^{\beta}(B_{d_{\Sigma}(\cdot)/2}(\cdot))}\leq C\rho^{\nu}\quad\text{in }\Sigma,$
where $C>0$ depends only on $n$, $\beta$, $\lambda_{i}$, and $\Sigma$.